OpenStudy (anonymous):

Solve x^2 + 10x + 24 = 0 by completing the square.

OpenStudy (anonymous):

This is where I'm stuck... x^2+10x+24=0 -24 -24 x^2+10x=-24 10/2 =5 ; 5^2=25 x^2+10x+25=-24 +25 x^2+10x+25=1 (x+5)(x+5)=1 (x+5)^2=1

OpenStudy (anonymous):

(x+5)^2 = 1 your next step is to get rid of the exponent on x+5. What's the opposite operation to "squaring" something?

OpenStudy (anonymous):

square root?

OpenStudy (anonymous):

bingo. So the square root of 1 equals x+5. What are the square roots of 1? There are 2 possibilities here

OpenStudy (anonymous):

1?

OpenStudy (anonymous):

and -1?

OpenStudy (anonymous):

yup! so you have x+5 = 1 and x+5 = -1 what are the 2 values x can be?

OpenStudy (anonymous):

x=-4 and x=-6?

OpenStudy (anonymous):

exactly. Good job!

OpenStudy (anonymous):

Thank you! :D